OpenStudy (katherinesmith):

write 1-3log5x as a single logarithm

OpenStudy (katherinesmith):

\[1 - 3\log _{5}x\]

OpenStudy (katherinesmith):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how far did you get

OpenStudy (katherinesmith):

nowhere.

OpenStudy (katherinesmith):

i don't know what a single logarithm looks like

jimthompson5910 (jim_thompson5910):

it looks like log(...) where the log is any base and there are no terms outside the log

OpenStudy (katherinesmith):

okay i know what you mean now

jimthompson5910 (jim_thompson5910):

try to get the given expression into the form I wrote

OpenStudy (katherinesmith):

i don't know how to go about doing that

OpenStudy (katherinesmith):

if someone could actually help me that would be great.

jimthompson5910 (jim_thompson5910):

hints: \[\large \log_{b}(b) = 1\] \[\large y\log_{b}(x) = \log_{b}(x^y)\] \[\large \log_{b}(x) - \log_{b}(y) = \log_{b}\left(\frac{x}{y}\right)\]

OpenStudy (katherinesmith):

okay i see your formulas but i have absolutely no idea what numbers are equivalent to what letters therefore i can't plug anything into a formula.

jimthompson5910 (jim_thompson5910):

you can replace b with 5 to go from \[\large \log_{b}(b) = 1\] to \[\large \log_{5}(5) = 1\]

OpenStudy (katherinesmith):

okay i see that.

OpenStudy (katherinesmith):

then what.

jimthompson5910 (jim_thompson5910):

use the other formulas I gave you

OpenStudy (katherinesmith):

what is y

jimthompson5910 (jim_thompson5910):

compare \[\large y\log_{b}(x) \] to \[\large 3\log_{5}(x)\]

OpenStudy (katherinesmith):

\[3\log _{5}(x) = \log _{5}(x ^{3})\]

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

we have this so far \[\large 1 - 3\log_{5}(x)\] \[\large \log_{5}(5) - 3\log_{5}(x)\] \[\large \log_{5}(5) - \log_{5}(x^3)\]

OpenStudy (katherinesmith):

\[\log _{5}(x) - \log _{5}(3) = \log _{5}(\frac{ x }{ 3 })\]

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

\[\large \log_{5}(5) - \log_{5}(x^3)\] turns into this \[\large \log_{5}\left(\frac{5}{x^3}\right) \] (see rule #3 given above)

jimthompson5910 (jim_thompson5910):

so this is the complete step by step picture

jimthompson5910 (jim_thompson5910):

\[\large 1 - 3\log_{5}(x)\] \[\large \log_{5}(5) - 3\log_{5}(x)\] \[\large \log_{5}(5) - \log_{5}(x^3)\] \[\large \log_{5}\left(\frac{5}{x^3}\right) \]

OpenStudy (katherinesmith):

and that's your answer huh

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (katherinesmith):

thank you a ton for your patience

jimthompson5910 (jim_thompson5910):

sure thing